1. Peggy-Sue loves to bake cookies and recently left her job at Cookie Monster, Inc., (making $125,000 per year) to start her own cookie-baking business. Her annual costs would include $40,000 for labor, $10,000 for rent, $35,000 for ingredients, and $5,000 for utilities. She withdrew $100,000 of her own savings to invest in the operation, which had been earning 10 percent interest per year. List Peggy-Sue’s total costs of operating her own cookie business, and specifically indicate which items are explicit costs and which are implicit costs. What are her total economic costs?
Answer #1

Explicit costs = labor+rent+ingredients+utilities = 40000+10000+35000+5000 = 90000

Implicit costs = forgone salary+forgone interest = 125000+100000*0.10 = 135000

Total economic costs = explicit+implicit costs = 90000+135000 = 225000